OpenStudy (anonymous):

A1=6 For n>1,An=A(n-1)+4

OpenStudy (anonymous):

This is an arithmetic series. What do you want to do with it?

OpenStudy (anonymous):

how to solve it

OpenStudy (anonymous):

solve for what?

OpenStudy (anonymous):

What are the first 4 terms of this sequence

OpenStudy (anonymous):

a1 = 6 a2 = 6 + 4 = 10 a3 = 10 + 4 = 14 a4 = 14 + 4 = 18